OpenStudy (anonymous):

Part I: Create two triangles that are not similar. Label your triangles and write the corresponding similarity statement. Describe, using complete sentences, the side lengths of the two triangles you created. Part II: Explain, using complete sentences, why these triangles are not similar. Note: The sides of the triangles don’t need to be measured; they can be made up as long as the two triangles are not similar. Also, the triangles don’t need to be submitted, just the triangle names and side lengths.

OpenStudy (ash2326):

It's easy, we need to make two triangles which are not similar Let the two triangles be ABC and XYZ let the side lengths be AB=4 BC=5 CA=6 XY=7 YZ=8 ZX=9 if you take the corresponding sides ratios \[\frac{AB}{XY}, \frac{BC}{YZ}, \frac{CA}{ZX}\] \[\frac{4}{7}\ne \frac 58 \ne \frac 6 9\] so the two triangles are not similar

OpenStudy (anonymous):

ty

OpenStudy (ash2326):